### v1

# Study of $\eta\rightarrow\pi^{+}\pi^{-}l^{+}l^{-}$

Study of $\eta\rightarrow\pi^{+}\pi^{-}l^{+}l^{-}$
M. Ablikim1, M. N. Achasov4,c, P. Adlarson76, O. Afedulidis3, X. C. Ai81, R. Aliberti35, A. Amoroso75A,75C, Q. An72,58,a, Y. Bai57, O. Bakina36, I. Balossino29A, Y. Ban46,h, H.-R. Bao64, V. Batozskaya1,44, K. Begzsuren32, N. Berger35, M. Berlowski44, M. Bertani28A, D. Bettoni29A, F. Bianchi75A,75C, E. Bianco75A,75C, A. Bortone75A,75C, I. Boyko36, R. A. Briere5, A. Brueggemann69, H. Cai77, X. Cai1,58, A. Calcaterra28A, G. F. Cao1,64, N. Cao1,64, S. A. Cetin62A, J. F. Chang1,58, G. R. Che43, G. Chelkov36,b, C. Chen43, C. H. Chen9, Chao Chen55, G. Chen1, H. S. Chen1,64, H. Y. Chen20, M. L. Chen1,58,64, S. J. Chen42, S. L. Chen45, S. M. Chen61, T. Chen1,64, X. R. Chen31,64, X. T. Chen1,64, Y. B. Chen1,58, Y. Q. Chen34, Z. J. Chen25,i, Z. Y. Chen1,64, S. K. Choi10A, G. Cibinetto29A, F. Cossio75C, J. J. Cui50, H. L. Dai1,58, J. P. Dai79, A. Dbeyssi18, R.  E. de Boer3, D. Dedovich36, C. Q. Deng73, Z. Y. Deng1, A. Denig35, I. Denysenko36, M. Destefanis75A,75C, F. De Mori75A,75C, B. Ding67,1, X. X. Ding46,h, Y. Ding40, Y. Ding34, J. Dong1,58, L. Y. Dong1,64, M. Y. Dong1,58,64, X. Dong77, M. C. Du1, S. X. Du81, Y. Y. Duan55, Z. H. Duan42, P. Egorov36,b, Y. H. Fan45, J. Fang1,58, J. Fang59, S. S. Fang1,64, W. X. Fang1, Y. Fang1, Y. Q. Fang1,58, R. Farinelli29A, L. Fava75B,75C, F. Feldbauer3, G. Felici28A, C. Q. Feng72,58, J. H. Feng59, Y. T. Feng72,58, M. Fritsch3, C. D. Fu1, J. L. Fu64, Y. W. Fu1,64, H. Gao64, X. B. Gao41, Y. N. Gao46,h, Yang Gao72,58, S. Garbolino75C, I. Garzia29A,29B, L. Ge81, P. T. Ge19, Z. W. Ge42, C. Geng59, E. M. Gersabeck68, A. Gilman70, K. Goetzen13, L. Gong40, W. X. Gong1,58, W. Gradl35, S. Gramigna29A,29B, M. Greco75A,75C, M. H. Gu1,58, Y. T. Gu15, C. Y. Guan1,64, A. Q. Guo31,64, L. B. Guo41, M. J. Guo50, R. P. Guo49, Y. P. Guo12,g, A. Guskov36,b, J. Gutierrez27, K. L. Han64, T. T. Han1, F. Hanisch3, X. Q. Hao19, F. A. Harris66, K. K. He55, K. L. He1,64, F. H. Heinsius3, C. H. Heinz35, Y. K. Heng1,58,64, C. Herold60, T. Holtmann3, P. C. Hong34, G. Y. Hou1,64, X. T. Hou1,64, Y. R. Hou64, Z. L. Hou1, B. Y. Hu59, H. M. Hu1,64, J. F. Hu56,j, S. L. Hu12,g, T. Hu1,58,64, Y. Hu1, G. S. Huang72,58, K. X. Huang59, L. Q. Huang31,64, X. T. Huang50, Y. P. Huang1, Y. S. Huang59, T. Hussain74, F. Hölzken3, N. Hüsken35, N. in der Wiesche69, J. Jackson27, S. Janchiv32, J. H. Jeong10A, Q. Ji1, Q. P. Ji19, W. Ji1,64, X. B. Ji1,64, X. L. Ji1,58, Y. Y. Ji50, X. Q. Jia50, Z. K. Jia72,58, D. Jiang1,64, H. B. Jiang77, P. C. Jiang46,h, S. S. Jiang39, T. J. Jiang16, X. S. Jiang1,58,64, Y. Jiang64, J. B. Jiao50, J. K. Jiao34, Z. Jiao23, S. Jin42, Y. Jin67, M. Q. Jing1,64, X. M. Jing64, T. Johansson76, S. Kabana33, N. Kalantar-Nayestanaki65, X. L. Kang9, X. S. Kang40, M. Kavatsyuk65, B. C. Ke81, V. Khachatryan27, A. Khoukaz69, R. Kiuchi1, O. B. Kolcu62A, B. Kopf3, M. Kuessner3, X. Kui1,64, N.  Kumar26, A. Kupsc44,76, W. Kühn37, J. J. Lane68, L. Lavezzi75A,75C, T. T. Lei72,58, Z. H. Lei72,58, M. Lellmann35, T. Lenz35, C. Li47, C. Li43, C. H. Li39, Cheng Li72,58, D. M. Li81, F. Li1,58, G. Li1, H. B. Li1,64, H. J. Li19, H. N. Li56,j, Hui Li43, J. R. Li61, J. S. Li59, K. Li1, L. J. Li1,64, L. K. Li1, Lei Li48, M. H. Li43, P. R. Li38,k,l, Q. M. Li1,64, Q. X. Li50, R. Li17,31, S. X. Li12, T.  Li50, W. D. Li1,64, W. G. Li1,a, X. Li1,64, X. H. Li72,58, X. L. Li50, X. Y. Li1,64, X. Z. Li59, Y. G. Li46,h, Z. J. Li59, Z. Y. Li79, C. Liang42, H. Liang72,58, H. Liang1,64, Y. F. Liang54, Y. T. Liang31,64, G. R. Liao14, Y. P. Liao1,64, J. Libby26, A.  Limphirat60, C. C. Lin55, D. X. Lin31,64, T. Lin1, B. J. Liu1, B. X. Liu77, C. Liu34, C. X. Liu1, F. Liu1, F. H. Liu53, Feng Liu6, G. M. Liu56,j, H. Liu38,k,l, H. B. Liu15, H. H. Liu1, H. M. Liu1,64, Huihui Liu21, J. B. Liu72,58, J. Y. Liu1,64, K. Liu38,k,l, K. Y. Liu40, Ke Liu22, L. Liu72,58, L. C. Liu43, Lu Liu43, M. H. Liu12,g, P. L. Liu1, Q. Liu64, S. B. Liu72,58, T. Liu12,g, W. K. Liu43, W. M. Liu72,58, X. Liu39, X. Liu38,k,l, Y. Liu81, Y. Liu38,k,l, Y. B. Liu43, Z. A. Liu1,58,64, Z. D. Liu9, Z. Q. Liu50, X. C. Lou1,58,64, F. X. Lu59, H. J. Lu23, J. G. Lu1,58, X. L. Lu1, Y. Lu7, Y. P. Lu1,58, Z. H. Lu1,64, C. L. Luo41, J. R. Luo59, M. X. Luo80, T. Luo12,g, X. L. Luo1,58, X. R. Lyu64, Y. F. Lyu43, F. C. Ma40, H. Ma79, H. L. Ma1, J. L. Ma1,64, L. L. Ma50, L. R. Ma67, M. M. Ma1,64, Q. M. Ma1, R. Q. Ma1,64, T. Ma72,58, X. T. Ma1,64, X. Y. Ma1,58, Y. Ma46,h, Y. M. Ma31, F. E. Maas18, M. Maggiora75A,75C, S. Malde70, Y. J. Mao46,h, Z. P. Mao1, S. Marcello75A,75C, Z. X. Meng67, J. G. Messchendorp13,65, G. Mezzadri29A, H. Miao1,64, T. J. Min42, R. E. Mitchell27, X. H. Mo1,58,64, B. Moses27, N. Yu. Muchnoi4,c, J. Muskalla35, Y. Nefedov36, F. Nerling18,e, L. S. Nie20, I. B. Nikolaev4,c, Z. Ning1,58, S. Nisar11,m, Q. L. Niu38,k,l, W. D. Niu55, Y. Niu 50, S. L. Olsen64, Q. Ouyang1,58,64, S. Pacetti28B,28C, X. Pan55, Y. Pan57, A.  Pathak34, Y. P. Pei72,58, M. Pelizaeus3, H. P. Peng72,58, Y. Y. Peng38,k,l, K. Peters13,e, J. L. Ping41, R. G. Ping1,64, S. Plura35, V. Prasad33, F. Z. Qi1, H. Qi72,58, H. R. Qi61, M. Qi42, T. Y. Qi12,g, S. Qian1,58, W. B. Qian64, C. F. Qiao64, X. K. Qiao81, J. J. Qin73, L. Q. Qin14, L. Y. Qin72,58, X. P. Qin12,g, X. S. Qin50, Z. H. Qin1,58, J. F. Qiu1, Z. H. Qu73, C. F. Redmer35, K. J. Ren39, A. Rivetti75C, M. Rolo75C, G. Rong1,64, Ch. Rosner18, S. N. Ruan43, N. Salone44, A. Sarantsev36,d, Y. Schelhaas35, K. Schoenning76, M. Scodeggio29A, K. Y. Shan12,g, W. Shan24, X. Y. Shan72,58, Z. J. Shang38,k,l, J. F. Shangguan16, L. G. Shao1,64, M. Shao72,58, C. P. Shen12,g, H. F. Shen1,8, W. H. Shen64, X. Y. Shen1,64, B. A. Shi64, H. Shi72,58, H. C. Shi72,58, J. L. Shi12,g, J. Y. Shi1, Q. Q. Shi55, S. Y. Shi73, X. Shi1,58, J. J. Song19, T. Z. Song59, W. M. Song34,1, Y.  J. Song12,g, Y. X. Song46,h,n, S. Sosio75A,75C, S. Spataro75A,75C, F. Stieler35, Y. J. Su64, G. B. Sun77, G. X. Sun1, H. Sun64, H. K. Sun1, J. F. Sun19, K. Sun61, L. Sun77, S. S. Sun1,64, T. Sun51,f, W. Y. Sun34, Y. Sun9, Y. J. Sun72,58, Y. Z. Sun1, Z. Q. Sun1,64, Z. T. Sun50, C. J. Tang54, G. Y. Tang1, J. Tang59, M. Tang72,58, Y. A. Tang77, L. Y. Tao73, Q. T. Tao25,i, M. Tat70, J. X. Teng72,58, V. Thoren76, W. H. Tian59, Y. Tian31,64, Z. F. Tian77, I. Uman62B, Y. Wan55, S. J. Wang 50, B. Wang1, B. L. Wang64, Bo Wang72,58, D. Y. Wang46,h, F. Wang73, H. J. Wang38,k,l, J. J. Wang77, J. P. Wang 50, K. Wang1,58, L. L. Wang1, M. Wang50, N. Y. Wang64, S. Wang12,g, S. Wang38,k,l, T.  Wang12,g, T. J. Wang43, W.  Wang73, W. Wang59, W. P. Wang35,72,o, W. P. Wang72,58, X. Wang46,h, X. F. Wang38,k,l, X. J. Wang39, X. L. Wang12,g, X. N. Wang1, Y. Wang61, Y. D. Wang45, Y. F. Wang1,58,64, Y. L. Wang19, Y. N. Wang45, Y. Q. Wang1, Yaqian Wang17, Yi Wang61, Z. Wang1,58, Z. L.  Wang73, Z. Y. Wang1,64, Ziyi Wang64, D. H. Wei14, F. Weidner69, S. P. Wen1, Y. R. Wen39, U. Wiedner3, G. Wilkinson70, M. Wolke76, L. Wollenberg3, C. Wu39, J. F. Wu1,8, L. H. Wu1, L. J. Wu1,64, X. Wu12,g, X. H. Wu34, Y. Wu72,58, Y. H. Wu55, Y. J. Wu31, Z. Wu1,58, L. Xia72,58, X. M. Xian39, B. H. Xiang1,64, T. Xiang46,h, D. Xiao38,k,l, G. Y. Xiao42, S. Y. Xiao1, Y.  L. Xiao12,g, Z. J. Xiao41, C. Xie42, X. H. Xie46,h, Y. Xie50, Y. G. Xie1,58, Y. H. Xie6, Z. P. Xie72,58, T. Y. Xing1,64, C. F. Xu1,64, C. J. Xu59, G. F. Xu1, H. Y. Xu67,2,p, M. Xu72,58, Q. J. Xu16, Q. N. Xu30, W. Xu1, W. L. Xu67, X. P. Xu55, Y. C. Xu78, Z. S. Xu64, F. Yan12,g, L. Yan12,g, W. B. Yan72,58, W. C. Yan81, X. Q. Yan1,64, H. J. Yang51,f, H. L. Yang34, H. X. Yang1, T. Yang1, Y. Yang12,g, Y. F. Yang1,64, Y. F. Yang43, Y. X. Yang1,64, Z. W. Yang38,k,l, Z. P. Yao50, M. Ye1,58, M. H. Ye8, J. H. Yin1, Junhao Yin43, Z. Y. You59, B. X. Yu1,58,64, C. X. Yu43, G. Yu1,64, J. S. Yu25,i, T. Yu73, X. D. Yu46,h, Y. C. Yu81, C. Z. Yuan1,64, J. Yuan45, J. Yuan34, L. Yuan2, S. C. Yuan1,64, Y. Yuan1,64, Z. Y. Yuan59, C. X. Yue39, A. A. Zafar74, F. R. Zeng50, S. H. Zeng63A,63B,63C,63D, X. Zeng12,g, Y. Zeng25,i, Y. J. Zeng59, Y. J. Zeng1,64, X. Y. Zhai34, Y. C. Zhai50, Y. H. Zhan59, A. Q. Zhang1,64, B. L. Zhang1,64, B. X. Zhang1, D. H. Zhang43, G. Y. Zhang19, H. Zhang81, H. Zhang72,58, H. C. Zhang1,58,64, H. H. Zhang59, H. H. Zhang34, H. Q. Zhang1,58,64, H. R. Zhang72,58, H. Y. Zhang1,58, J. Zhang81, J. Zhang59, J. J. Zhang52, J. L. Zhang20, J. Q. Zhang41, J. S. Zhang12,g, J. W. Zhang1,58,64, J. X. Zhang38,k,l, J. Y. Zhang1, J. Z. Zhang1,64, Jianyu Zhang64, L. M. Zhang61, Lei Zhang42, P. Zhang1,64, Q. Y. Zhang34, R. Y. Zhang38,k,l, S. H. Zhang1,64, Shulei Zhang25,i, X. D. Zhang45, X. M. Zhang1, X. Y. Zhang50, Y.  Zhang73, Y. Zhang1, Y.  T. Zhang81, Y. H. Zhang1,58, Y. M. Zhang39, Yan Zhang72,58, Z. D. Zhang1, Z. H. Zhang1, Z. L. Zhang34, Z. Y. Zhang43, Z. Y. Zhang77, Z. Z.  Zhang45, G. Zhao1, J. Y. Zhao1,64, J. Z. Zhao1,58, L. Zhao1, Lei Zhao72,58, M. G. Zhao43, N. Zhao79, R. P. Zhao64, S. J. Zhao81, Y. B. Zhao1,58, Y. X. Zhao31,64, Z. G. Zhao72,58, A. Zhemchugov36,b, B. Zheng73, B. M. Zheng34, J. P. Zheng1,58, W. J. Zheng1,64, Y. H. Zheng64, B. Zhong41, X. Zhong59, H.  Zhou50, J. Y. Zhou34, L. P. Zhou1,64, S.  Zhou6, X. Zhou77, X. K. Zhou6, X. R. Zhou72,58, X. Y. Zhou39, Y. Z. Zhou12,g, A. N. Zhu64, J. Zhu43, K. Zhu1, K. J. Zhu1,58,64, K. S. Zhu12,g, L. Zhu34, L. X. Zhu64, S. H. Zhu71, T. J. Zhu12,g, W. D. Zhu41, Y. C. Zhu72,58, Z. A. Zhu1,64, J. H. Zou1, J. Zu72,58 (BESIII Collaboration) 1 Institute of High Energy Physics, Beijing 100049, People’s Republic of China
2 Beihang University, Beijing 100191, People’s Republic of China
3 Bochum Ruhr-University, D-44780 Bochum, Germany
4 Budker Institute of Nuclear Physics SB RAS (BINP), Novosibirsk 630090, Russia
5 Carnegie Mellon University, Pittsburgh, Pennsylvania 15213, USA
6 Central China Normal University, Wuhan 430079, People’s Republic of China
7 Central South University, Changsha 410083, People’s Republic of China
8 China Center of Advanced Science and Technology, Beijing 100190, People’s Republic of China
9 China University of Geosciences, Wuhan 430074, People’s Republic of China
10 Chung-Ang University, Seoul, 06974, Republic of Korea
11 COMSATS University Islamabad, Lahore Campus, Defence Road, Off Raiwind Road, 54000 Lahore, Pakistan
12 Fudan University, Shanghai 200433, People’s Republic of China
13 GSI Helmholtzcentre for Heavy Ion Research GmbH, D-64291 Darmstadt, Germany
14 Guangxi Normal University, Guilin 541004, People’s Republic of China
15 Guangxi University, Nanning 530004, People’s Republic of China
16 Hangzhou Normal University, Hangzhou 310036, People’s Republic of China
17 Hebei University, Baoding 071002, People’s Republic of China
18 Helmholtz Institute Mainz, Staudinger Weg 18, D-55099 Mainz, Germany
19 Henan Normal University, Xinxiang 453007, People’s Republic of China
20 Henan University, Kaifeng 475004, People’s Republic of China
21 Henan University of Science and Technology, Luoyang 471003, People’s Republic of China
22 Henan University of Technology, Zhengzhou 450001, People’s Republic of China
23 Huangshan College, Huangshan 245000, People’s Republic of China
24 Hunan Normal University, Changsha 410081, People’s Republic of China
25 Hunan University, Changsha 410082, People’s Republic of China
26 Indian Institute of Technology Madras, Chennai 600036, India
27 Indiana University, Bloomington, Indiana 47405, USA
28 INFN Laboratori Nazionali di Frascati , (A)INFN Laboratori Nazionali di Frascati, I-00044, Frascati, Italy; (B)INFN Sezione di Perugia, I-06100, Perugia, Italy; (C)University of Perugia, I-06100, Perugia, Italy
29 INFN Sezione di Ferrara, (A)INFN Sezione di Ferrara, I-44122, Ferrara, Italy; (B)University of Ferrara, I-44122, Ferrara, Italy
30 Inner Mongolia University, Hohhot 010021, People’s Republic of China
31 Institute of Modern Physics, Lanzhou 730000, People’s Republic of China
32 Institute of Physics and Technology, Peace Avenue 54B, Ulaanbaatar 13330, Mongolia
33 Instituto de Alta Investigación, Universidad de Tarapacá, Casilla 7D, Arica 1000000, Chile
34 Jilin University, Changchun 130012, People’s Republic of China
35 Johannes Gutenberg University of Mainz, Johann-Joachim-Becher-Weg 45, D-55099 Mainz, Germany
36 Joint Institute for Nuclear Research, 141980 Dubna, Moscow region, Russia
37 Justus-Liebig-Universitaet Giessen, II. Physikalisches Institut, Heinrich-Buff-Ring 16, D-35392 Giessen, Germany
38 Lanzhou University, Lanzhou 730000, People’s Republic of China
39 Liaoning Normal University, Dalian 116029, People’s Republic of China
40 Liaoning University, Shenyang 110036, People’s Republic of China
41 Nanjing Normal University, Nanjing 210023, People’s Republic of China
42 Nanjing University, Nanjing 210093, People’s Republic of China
43 Nankai University, Tianjin 300071, People’s Republic of China
44 National Centre for Nuclear Research, Warsaw 02-093, Poland
45 North China Electric Power University, Beijing 102206, People’s Republic of China
46 Peking University, Beijing 100871, People’s Republic of China
47 Qufu Normal University, Qufu 273165, People’s Republic of China
48 Renmin University of China, Beijing 100872, People’s Republic of China
49 Shandong Normal University, Jinan 250014, People’s Republic of China
50 Shandong University, Jinan 250100, People’s Republic of China
51 Shanghai Jiao Tong University, Shanghai 200240, People’s Republic of China
52 Shanxi Normal University, Linfen 041004, People’s Republic of China
53 Shanxi University, Taiyuan 030006, People’s Republic of China
54 Sichuan University, Chengdu 610064, People’s Republic of China
55 Soochow University, Suzhou 215006, People’s Republic of China
56 South China Normal University, Guangzhou 510006, People’s Republic of China
57 Southeast University, Nanjing 211100, People’s Republic of China
58 State Key Laboratory of Particle Detection and Electronics, Beijing 100049, Hefei 230026, People’s Republic of China
59 Sun Yat-Sen University, Guangzhou 510275, People’s Republic of China
60 Suranaree University of Technology, University Avenue 111, Nakhon Ratchasima 30000, Thailand
61 Tsinghua University, Beijing 100084, People’s Republic of China
62 Turkish Accelerator Center Particle Factory Group, (A)Istinye University, 34010, Istanbul, Turkey; (B)Near East University, Nicosia, North Cyprus, 99138, Mersin 10, Turkey
63 University of Bristol, (A)H H Wills Physics Laboratory; (B)Tyndall Avenue; (C)Bristol; (D)BS8 1TL
64 University of Chinese Academy of Sciences, Beijing 100049, People’s Republic of China
65 University of Groningen, NL-9747 AA Groningen, The Netherlands
66 University of Hawaii, Honolulu, Hawaii 96822, USA
67 University of Jinan, Jinan 250022, People’s Republic of China
68 University of Manchester, Oxford Road, Manchester, M13 9PL, United Kingdom
69 University of Muenster, Wilhelm-Klemm-Strasse 9, 48149 Muenster, Germany
70 University of Oxford, Keble Road, Oxford OX13RH, United Kingdom
71 University of Science and Technology Liaoning, Anshan 114051, People’s Republic of China
72 University of Science and Technology of China, Hefei 230026, People’s Republic of China
73 University of South China, Hengyang 421001, People’s Republic of China
74 University of the Punjab, Lahore-54590, Pakistan
75 University of Turin and INFN, (A)University of Turin, I-10125, Turin, Italy; (B)University of Eastern Piedmont, I-15121, Alessandria, Italy; (C)INFN, I-10125, Turin, Italy
76 Uppsala University, Box 516, SE-75120 Uppsala, Sweden
77 Wuhan University, Wuhan 430072, People’s Republic of China
78 Yantai University, Yantai 264005, People’s Republic of China
79 Yunnan University, Kunming 650500, People’s Republic of China
80 Zhejiang University, Hangzhou 310027, People’s Republic of China
81 Zhengzhou University, Zhengzhou 450001, People’s Republic of China
a Deceased
b Also at the Moscow Institute of Physics and Technology, Moscow 141700, Russia
c Also at the Novosibirsk State University, Novosibirsk, 630090, Russia
d Also at the NRC ”Kurchatov Institute”, PNPI, 188300, Gatchina, Russia
e Also at Goethe University Frankfurt, 60323 Frankfurt am Main, Germany
f Also at Key Laboratory for Particle Physics, Astrophysics and Cosmology, Ministry of Education; Shanghai Key Laboratory for Particle Physics and Cosmology; Institute of Nuclear and Particle Physics, Shanghai 200240, People’s Republic of China
g Also at Key Laboratory of Nuclear Physics and Ion-beam Application (MOE) and Institute of Modern Physics, Fudan University, Shanghai 200443, People’s Republic of China
h Also at State Key Laboratory of Nuclear Physics and Technology, Peking University, Beijing 100871, People’s Republic of China
i Also at School of Physics and Electronics, Hunan University, Changsha 410082, China
j Also at Guangdong Provincial Key Laboratory of Nuclear Science, Institute of Quantum Matter, South China Normal University, Guangzhou 510006, China
k Also at MOE Frontiers Science Center for Rare Isotopes, Lanzhou University, Lanzhou 730000, People’s Republic of China
l Also at Lanzhou Center for Theoretical Physics, Lanzhou University, Lanzhou 730000, People’s Republic of China
m Also at the Department of Mathematical Sciences, IBA, Karachi 75270, Pakistan
n Also at Ecole Polytechnique Federale de Lausanne (EPFL), CH-1015 Lausanne, Switzerland
o Also at Helmholtz Institute Mainz, Staudinger Weg 18, D-55099 Mainz, Germany
p Also at School of Physics, Beihang University, Beijing 100191 , China
Abstract
Using a sample of $(10087\pm 44)\times 10^{6}$ $J/\psi$ events accumulated with the BESIII detector, we analyze the decays $\eta\rightarrow\pi^{+}\pi^{-}l^{+}l^{-}$ ($l=e$ or $\mu$) via the process $J/\psi\rightarrow\gamma\eta$. The branching fraction of $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ is measured to be $\mathcal{B}(\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-})=(3.07\pm 0.12_{\rm{stat.}}\pm 0.19_{\rm{syst.}})\times 10^{-4}$. No signal events are observed for the $\eta\rightarrow\pi^{+}\pi^{-}\mu^{+}\mu^{-}$ decay, leading to an upper limit on the branching fraction of $\mathcal{B}(\eta\rightarrow\pi^{+}\pi^{-}\mu^{+}\mu^{-})<4.0\times 10^{-7}$ at the 90% confidence level. Furthermore, the $CP$-violation asymmetry parameter is found to be $\mathcal{A}_{CP}(\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-})=(-4.04\pm 4.69_{\rm{stat.}}\pm 0.14_{\rm{syst.}})\%$, showing no evidence of $CP$-violation with current statistics. Additionally, we extract the transition form factor from the decay amplitude of $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$. Finally, axion-like particles are searched for via the decay $\eta\rightarrow\pi^{+}\pi^{-}a,a\rightarrow e^{+}e^{-}$, and upper limits on this branching fraction relative to that of $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ are presented as a function of the axion-like particle mass in the range $5-200\ \mathrm{MeV}/c^{2}$.
I INTRODUCTION
The decays $\eta\rightarrow\pi^{+}\pi^{-}l^{+}l^{-}$ ($l=e$ or $\mu$) involve contributions from the box anomaly of quantum chromodynamics (QCD) and have been studied using various models, such as the hidden gauge Borasoy and Nissler (2007), the chiral unitary approach (Unitary $\chi$PT) Borasoy and Nissler (2007), and the vector meson dominance (VMD) Petri (2010) models. Table 1 displays the theoretical predictions and experimental results for the branching fractions ($\mathcal{B}$) of the $\eta\rightarrow\pi^{+}\pi^{-}l^{+}l^{-}$ decays. Electromagnetic decays of mesons depend on the coupling of photons to the electric charge distribution of the quark fields. This electromagnetic meson structure is described by the transition form factor (TFF), which is a function of the momentum transfer. More importantly, the TFF determines the size of certain hadronic quantum corrections in the anomalous magnetic moment of the muon, $(g-2)_{\mu}$ Danilkin et al. (2019).
Table 1: Different theoretical predictions and previous experimental measurements of $\mathcal{B}(\eta\rightarrow\pi^{+}\pi^{-}l^{+}l^{-})$.
$\mathcal{B}(\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-})$ $(10^{-4})$ $\mathcal{B}(\eta\rightarrow\pi^{+}\pi^{-}\mu^{+}\mu^{-})$ $(10^{-9})$
Unitary $\chi$PT Borasoy and Nissler (2007) $2.99^{+0.06}_{-0.09}$ $7.50^{+1.80}_{-0.70}$
Hidden gauge Petri (2010) $3.14\pm 0.17$ $8.65\pm 0.39$
VMD Petri (2010) $3.02\pm 0.12$ $8.64\pm 0.25$
CMD-2 Akhmetshin et al. (2001) $3.7^{+2.5}_{-1.8}\pm 3.0$..
WASA Berlowski et al. (2008) $4.3^{+0.2}_{-1.6}\pm 0.4$ $<3.6\times 10^{5}$
KLOE Ambrosino et al. (2009) $2.68\pm 0.09\pm 0.07$..
In the year 2000, the KTeV experiment observed a significant $CP$ asymmetry in the distribution of the T-odd angle $\varphi$ in $K^{0}_{L}\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ decay Alavi-Harati et al. (2000), where $\varphi$ is the angle between the $e^{+}e^{-}$ decay plane and the $\pi^{+}\pi^{-}$ decay plane in the $K^{0}_{L}$ center-of-mass system. However, so far there is no experimental indication of $CP$-violation in flavor-conserving reactions. The $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ decay provides an opportunity to search for an analogous $CP$-violation asymmetry in the angle distribution between the $e^{+}e^{-}$ and $\pi^{+}\pi^{-}$ decay planes. This has been explored in $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ decay in the KLOE experiment, and the $CP$-violation asymmetry is determined to be $\mathcal{A}_{CP}=(-0.6\pm 2.5\pm 1.8)\%$ Ambrosino et al. (2009), which is consistent with zero.
The hadronic decay channels of the $\eta$ meson could also provide signals of the QCD axion, a dark photon or other axion-like particles (ALPs) Landini and Meggiolaro (2020); Alves and Gonzàlez-Solís (2024) that couple with hadrons. Hints of a new light bosonic state with mass around $17\ \mathrm{MeV/c^{2}}$ Krasznahorkay et al. (2016, 2019) were observed by the ATOMKI Collaboration via measuring the angle between $e^{+}e^{-}$ pairs Buttazzo et al. (2021); Ellwanger and Moretti (2016); Liu et al. (2021), and a light pseudoscalar particle $a$ decaying to $e^{+}e^{-}$ was proposed to explain the ATOMKI anomaly. The ALPs could also cause a deviation from the expected value of the electron anomalous magnetic moment Parker et al. (2018); Morel et al. (2020); Andreev et al. (2021).
The ten billion $J/\psi$ events collected with the BESIII detector during 2009-2019 Ablikim et al. (2022) offer an excellent opportunity to measure the TFF, and to search for $CP$-violation and hadronically coupled ALPs.
II BESIII DETECTOR
The BESIII detector Ablikim et al. (2010) records symmetric $e^{+}e^{-}$ collisions provided by the BEPCII storage ring Yu et al. (2016) in the center-of-mass energy range from 2.0 to 4.95 GeV, with a peak luminosity of $1\times 10^{33}\;\text{cm}^{-2}\text{s}^{-1}$ achieved at $\sqrt{s}=3.77\;\text{GeV}$. BESIII has collected large data samples in this energy region Ablikim et al. (2020). The cylindrical core of the BESIII detector covers 93% of the full solid angle and consists of a helium-based multilayer drift chamber (MDC), a plastic scintillator time-of-flight system (TOF), and a CsI(Tl) electromagnetic calorimeter (EMC), which are all enclosed in a superconducting solenoidal magnet providing a 1.0 T magnetic field. The magnetic field was 0.9 T in 2012, which affects 10% of the total $J/\psi$ data. The solenoid is supported by an octagonal flux-return yoke with resistive plate counter muon identification modules interleaved with steel. The charged-particle momentum resolution at $1~{}{\rm GeV}/c$ is $0.5\%$, and the ${\rm d}E/{\rm d}x$ resolution is $6\%$ for electrons from Bhabha scattering. The EMC measures photon energies with a resolution of $2.5\%$ ($5\%$) at $1$ GeV in the barrel (end cap) region. The time resolution in the TOF barrel region is 68 ps, while that in the end cap region was 110 ps. The end cap TOF system was upgraded in 2015 using multigap resistive plate chamber technology, providing a time resolution of 60 ps Cao et al. (2020). About 13% of the data used here benefits from this upgrade.
III DATA SAMPLE AND MONTE CARLO SIMULATION
This analysis is based on $(10087\pm 44)\times 10^{6}$ $J/\psi$ events collected with the BESIII detector from 2009 to 2019 Ablikim et al. (2022); it uses the radiative decay $J/\psi\rightarrow\gamma\eta$, resulting in a sample of about $1.1\times 10^{7}$ $\eta$ events.
The estimation of backgrounds and the determination of detection efficiencies are performed with Monte Carlo (MC) simulations. The BESIII detectors are modeled with GEANT4 Agostinelli et al. (2003); Allison et al. (2006); Huang et al. (2022). The production of the $J/\psi$ resonance is implemented with the MC event generator KKMC Jadach et al. (2000, 2001), while the decays are simulated by EVTGEN Ping (2008); Lange (2001). The possible backgrounds are studied using a sample of $J/\psi$ inclusive events in which the known decays of $J/\psi$ are modeled with branching fractions set to the world average values in the PDG Navas et al. (2024), while the unknown decays are generated with the LUNDCHARM model Chen et al. (2000). For this analysis, specific generators for the $\eta\rightarrow\pi^{+}\pi^{-}l^{+}l^{-}$ Zhang et al. (2012), $\eta\rightarrow\pi^{+}\pi^{-}\pi^{0}$ Qin et al. (2018), $\eta\rightarrow\gamma\pi^{+}\pi^{-}$ Qin et al. (2018), $\eta\rightarrow\gamma e^{+}e^{-}$ Qin et al. (2018), $\eta^{\prime}\rightarrow\eta\pi^{+}\pi^{-}$ Qin et al. (2018) decays are developed based on theoretical amplitudes.
IV Event selection and background analysis
The final state of interest is studied through the decay chain $J/\psi\rightarrow\gamma\eta$, $\eta\rightarrow\pi^{+}\pi^{-}l^{+}l^{-}$. Each event is required to contain at least one good photon candidate, and four charged track candidates with a total charge of zero. Charged tracks detected in the MDC are required to be within a polar angle $(\theta)$ range of $|\cos\theta|\leq 0.93$, where $\theta$ is defined with respect to the $z$-axis, which is the symmetry axis of the MDC. The distance of closest approach to the interaction point (IP) must be less than 10 cm along the $z$-axis and less than 1 cm in the transverse plane. Photon candidates are identified using showers in the EMC. The deposited energy of each shower must be larger than 25 MeV in the barrel region ($|\cos\theta|<0.8$) or larger than 50 MeV in the end cap region ($(0.86<|\cos\theta|<0.92)$). To exclude showers that originate from charged tracks, the angle subtended by the EMC shower and the position of the closest charged track at the EMC must be greater than 15 degrees as measured from the IP. To suppress electronic noise and showers unrelated to the event, the difference between the EMC time and the event start time is required to be within [0, 700] ns.
[…]
For the decay $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$, a requirement of $\chi^{2}_{\rm{sum}}(\pi^{+}\pi^{-}e^{+}e^{-})<50$ is imposed. The primary peaking background comes from the $\eta\rightarrow\gamma\pi^{+}\pi^{-}$ events, where the photon converts to an $e^{+}e^{-}$ pair. To exclude these background events, a photon conversion finder algorithm Xu and He (2012) is applied to the selected $e^{+}e^{-}$ pair. The distribution of $R_{xy}$, defined as the projected distance from the $e^{+}e^{-}$-vertex position to the IP in the $x-y$ plane Xu and He (2012), is shown in Fig. 1.
Photon conversion events are identified based on their proximity to the beam pipe ($R_{xy}\approx 3.5$ cm) and the inner wall of the MDC ($R_{xy}\approx 6.5$ cm). The observable $\Phi_{ee}$ is the lab-frame opening angle of the $e^{+}e^{-}$ pair Adare et al. (2010), and the two-dimensional distribution of $\Phi_{ee}$ vs. $R_{xy}$ is shown in Fig. 2. For photon conversion events, $\Phi_{ee}$ is expected to be close to zero. After the application of the photon conversion veto $\Phi_{ee}<\frac{\pi}{3}$ when $2.0\ \mathrm{cm}<R_{xy}<7.5\ \mathrm{cm}$, most conversion events are rejected. The remaining background events from the inclusive MC sample are studied, and the dominant background channels listed in Table 2. Dedicated exclusive MC samples are then generated to better estimate their contributions.
Figure 1: The distribution of $R_{xy}$. The dots with error bars represent the data, the red dashed histogram is the MC signal shape, the green dashed histogram is the $J/\psi\rightarrow\gamma\eta,\eta\rightarrow\gamma\pi^{+}\pi^{-}$ MC shape, and the blue solid histogram is the sum of the MC signal and MC background from $J/\psi\rightarrow\gamma\eta,\eta\rightarrow\gamma\pi^{+}\pi^{-}$.
(a)Refer to caption(b)
Figure 2: The photon conversion veto criterion in the $\Phi_{ee}$ vs. $R_{xy}$ plane, shown with events from (a) data and (b) $\eta\rightarrow\gamma\pi^{+}\pi^{-}$ background MC sample. Events inside the red rectangle are rejected.
Table 2: Main background processes and normalized yields of the $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ decay. For the $\mathcal{B}$ determination, “--” indicates that the number of background events are float in the fit. “Other” refers to the analyses for TFF, $\mathcal{A}_{CP}$, and an ALP. A more detailed explanation is provided in the subsequent sections.
Background mode Normalized yields
$\mathcal{B}$ Other
$J/\psi\rightarrow\gamma\eta^{\prime},\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\eta,\eta\rightarrow\gamma e^{+}e^{-}$ $85\pm 6$ $8\pm 2$
$J/\psi\rightarrow\gamma\eta,\eta\rightarrow\pi^{+}\pi^{-}\pi^{0}$ $312\pm 13$ $18\pm 2$
$J/\psi\rightarrow\gamma\eta,\eta\rightarrow\gamma\pi^{+}\pi^{-}$ $42\pm 3$ $29\pm 3$
$J/\psi\rightarrow\gamma\pi^{+}\pi^{-}\pi^{+}\pi^{-}$- $16\pm 2$
For the decay $\eta\rightarrow\pi^{+}\pi^{-}\mu^{+}\mu^{-}$, a requirement of $\chi^{2}_{\rm{sum}}(\pi^{+}\pi^{-}\mu^{+}\mu^{-})<50$ is imposed. The signal window is determined by fitting the $\pi^{+}\pi^{-}\mu^{+}\mu^{-}$ invariant mass spectrum of the signal MC sample. The signal shape is represented by a double Gaussian function, and the signal window is defined as $[\mu-3\sigma,\mu+3\sigma]$, where $\mu$ and $\sigma$ are the weighted mean and standard deviation from the fit. In this case, the signal window is set to $0.531\ \mathrm{GeV}/c^{2}<M(\pi^{+}\pi^{-}\mu^{+}\mu^{-})<0.567\ \mathrm{GeV}/c^{2}$, and there are no events in this $\eta$ mass region, as displayed in Fig. 3.
Figure 3: The $\pi^{+}\pi^{-}\mu^{+}\mu^{-}$ invariant mass distribution of selected candidates in data.
V Analysis of $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$
V.1 Branching Fraction Measurement
To extract the number of $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ events, an unbinned maximum likelihood fit is performed to the $\pi^{+}\pi^{-}e^{+}e^{-}$ invariant mass spectrum. The signal and all the background shapes are taken from the MC simulations. The number of background events from $J/\psi\rightarrow\gamma\pi^{+}\pi^{-}\pi^{+}\pi^{-}$ is determined from the fit, and all other background events are fixed according to the branching fractions from the PDG Navas et al. (2024), as summarized in Table 2. The fit yields $N_{\rm{sig}}=680\pm 27$ signal events, and the result of the invariant mass of $\pi^{+}\pi^{-}e^{+}e^{-}$ is shown in Fig. 4. The fit $\chi^{2}$ per degree of freedom ($ndf$) is $\chi^{2}/ndf=75.1/98$. The branching fraction is determined by
$\mathcal{B}(\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-})$ $=\frac{N_{\rm{sig}}}{N_{J/\psi}\,\mathcal{B}(J/\psi\rightarrow\gamma\eta)\,\varepsilon}.$ (1)
Here $N_{J/\psi}$ is the number of $J/\psi$ events Ablikim et al. (2022) and $\mathcal{B}(J/\psi\rightarrow\gamma\eta)$ is the branching fraction of $J/\psi\rightarrow\gamma\eta$ Navas et al. (2024). With a detection efficiency of $\varepsilon=(20.12\pm 0.04)\%$, the branching fraction of $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ is calculated to be $(3.07\pm 0.12)\times 10^{-4}$, where the uncertainty is statistical only.
Figure 4: Fit to the invariant mass distribution of $\pi^{+}\pi^{-}e^{+}e^{-}$. The dots with error bars represent the data, the red solid line is the MC signal shape, and the blue solid line is the total fit result. The blue dot-dash line (Background I) is the $J/\psi\rightarrow\gamma\eta^{\prime},\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\eta,\eta\rightarrow\gamma e^{+}e^{-}$ MC shape. The green long-dash line (Background II) is the $J/\psi\rightarrow\gamma\eta,\eta\rightarrow\pi^{+}\pi^{-}\pi^{0}$ MC shape. The purple dashed line (Background III) is the $J/\psi\rightarrow\gamma\eta,\eta\rightarrow\gamma\pi^{+}\pi^{-}$ MC shape. The pink short-dash line (Background IV) is the $J/\psi\rightarrow\gamma\pi^{+}\pi^{-}\pi^{+}\pi^{-}$ MC shape.
V.2 Form Factor Measurement
The form factor is written as
$\mathcal{M}(s_{\pi\pi},s_{ee})=\mathcal{M}_{\eta}\times\mathrm{VMD}(s_{\pi\pi},s_{ee}).$ (2)
Here $\mathcal{M}_{\eta}$ includes factors related to decay constants and pseudoscalar-octet mixing, and $\mathrm{VMD}(s_{\pi\pi},s_{ee})$ is the VMD factor which included propagator effects. The four-momenta of the $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ decay is $P_{\eta}=P_{\pi^{+}}+P_{\pi^{-}}+P_{e^{+}}+P_{e^{-}}$, and the VMD function arguments are Gao (2002)
$s_{\pi\pi}=(P_{\pi^{+}}+P_{\pi^{-}})^{2},s_{ee}=(P_{e^{+}}+P_{e^{-}})^{2}.$ (3)
The VMD factor is derived from the VMD model with finite-width corrections Gounaris and Sakurai (1968) as
$\mathrm{VMD}(s_{\pi\pi},s_{ee})=$ $1-\frac{3}{4}(c_{1}-c_{2}+c_{3})$ (4)
$+\frac{3}{4}(c_{1}-c_{2}-c_{3})\frac{m^{2}_{V}}{m^{2}_{V}-s_{ee}-im_{V}\Gamma(s_{ee})}$
$+\frac{3}{2}c_{3}\frac{m^{2}_{V}}{m^{2}_{V}-s_{\pi\pi}-im^{2}_{V}\Gamma(s_{\pi\pi})}$
$\cdot\frac{m^{2}_{V}}{m^{2}_{V}-s_{ee}-im_{V}\Gamma(s_{ee})}.$
Here $m_{V}$ is the mass of the vector meson and $\Gamma(s)$ is its total width Zhang et al. (2012)
$\Gamma(s)\,=\,\left(\frac{\Gamma_{\rho(770)}\,\sqrt{s}}{m_{V}}\right)\left(\frac{1-\frac{4m^{2}_{i}}{s}}{1-\frac{4m^{2}_{i}}{m^{2}_{V}}}\right)^{\frac{3}{2}}\Theta(s-4m^{2}_{i}),$ (5)
where $i$ is $e\ \mathrm{or}\ \pi$, $\Gamma_{\rho(770)}=149.1$ MeV is the width of $\rho(770)$, and $\Theta$ is the Heaviside step function. By adjusting the values of the $c_{i}$-parameters Benayoun et al. (2010), one can switch between the various VMD models: (I) hidden gauge model ($c_{1}-c_{2}=c_{3}=1$); (II) full VMD model ($c_{1}-c_{2}=1/3,c_{3}=1$);
To extract the parameter $m_{V}$ in the VMD factor, an unbinned maximum likelihood fit to the $e^{+}e^{-}$ and $\pi^{+}\pi^{-}$ invariant mass spectra in data is performed using MINUIT James and Roos (1975). In the fit, $M(\pi^{+}\pi^{-}e^{+}e^{-})$ is required to be in the $\eta$ mass region of $(0.53-0.57)\mathrm{GeV}/c^{2}$. The backgrounds are generated by MC simulation with specific generators as described in Sec. III. The number of background events of $J/\psi\rightarrow\gamma\pi^{+}\pi^{-}\pi^{+}\pi^{-}$ is obtained by fitting the $\pi^{+}\pi^{-}e^{+}e^{-}$ invariant mass spectrum, and other background yields, listed in Table 2, are estimated using PDG branching fractions Navas et al. (2024). The probability of observing the $i^{th}$ event characterized by the measured four-momenta $\xi_{i}$ of final-state particles is
$\mathcal{P}(\xi_{i})=\frac{|\mathcal{A}(\xi_{i})|^{2}\,\varepsilon(\xi_{i})}{\int|\mathcal{A}(\xi)|^{2}\,\varepsilon(\xi)\,\mathrm{d}\xi},$ (6)
where $\mathcal{A}$ is the amplitude and $\varepsilon(\xi_{i})$ is the detection efficiency.
The normalization factor in the denominator is computed by integrating over all kinematic variables. The fit is done by minimizing the negative log-likelihood value
$-\ln{\mathcal{L}}=-\omega^{\prime}[$ $\sum^{N_{\mathrm{data}}}_{i=1}\ln{\mathcal{P}(\xi_{i})}-\omega_{\mathrm{bkg1}}\sum^{N_{\mathrm{bkg1}}}_{j=1}\ln{\mathcal{P}(\xi_{j})}$ (7)
$-\omega_{\mathrm{bkg2}}\sum^{N_{\mathrm{bkg2}}}_{k=1}\ln{\mathcal{P}(\xi_{k})}-...\,],$
where $i$ runs over all accepted events, and $j$, $k$, … run over the other considered background events. Their corresponding numbers of events are denoted by $N_{\mathrm{data}}$, $N_{\mathrm{bkg1}}$ and $N_{\mathrm{bkg2}}$. $\omega_{\mathrm{bkg1}}=\frac{N^{\prime}_{\mathrm{bkg1}}}{N_{\mathrm{bkg1}}}$ and $\omega_{\mathrm{bkg2}}=\frac{N^{\prime}_{\mathrm{bkg2}}}{N_{\mathrm{bkg2}}}$ are the weights of the backgrounds, where $N^{\prime}_{\mathrm{bkg1}}$ and $N^{\prime}_{\mathrm{bkg2}}$ are the contributions according to individual branching fractions taken from the PDG. To obtain an unbiased uncertainty estimation, the normalization factor $\omega^{\prime}$ Langenbruch (2022) used is
$\omega^{\prime}=\frac{N_{\mathrm{data}}-N_{\mathrm{bkg1}}\omega_{\mathrm{bkg1}}-N_{\mathrm{bkg2}}\omega_{\mathrm{bkg2}}+...}{N_{\mathrm{data}}+N_{\mathrm{bkg1}}\omega^{2}_{\mathrm{bkg1}}+N_{\mathrm{bkg2}}\omega^{2}_{\mathrm{bkg2}}+...}.$ (8)
The fit results of Model I and Model II are both in good agreement with data; Fig. 5 shows the $\pi^{+}\pi^{-}$ and $e^{+}e^{-}$ invariant mass spectrum fits. The results for $m_{V}$ in Model I and Model II are $m_{V}=749\pm 54\ \mathrm{MeV}/c^{2}$ and $m_{V}=748\pm 53\ \mathrm{MeV}/c^{2}$, respectively. The $\chi^{2}/ndf(\pi^{+}\pi^{-},e^{+}e^{-})$ are calculated to be (I) $73.6/41,22.6/36$ and (II) $72.9/41,22.6/36$. A larger $\eta$ data sample, potentially obtainable from experiments like the Super $\tau-$Charm Facility Achasov et al. (2024), is necessary to achieve a more precise measurement.
(a)Refer to caption(b)
Figure 5: Fit to the invariant mass distribution of the (a) $e^{+}e^{-}$ and (b) $\pi^{+}\pi^{-}$ pairs for the $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ decay mode. The points with error bars represent the data. The red and blue solid histogram is the total fit result. The light blue dashed histogram (Background I) is the $J/\psi\rightarrow\gamma\eta^{\prime},\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\eta,\eta\rightarrow\gamma e^{+}e^{-}$ MC shape. The green dashed histogram (Background II) is the $J/\psi\rightarrow\gamma\eta,\eta\rightarrow\pi^{+}\pi^{-}\pi^{0}$ MC shape. The purple dashed histogram (Background III) is the $J/\psi\rightarrow\gamma\eta,\eta\rightarrow\gamma\pi^{+}\pi^{-}$ MC shape. The pink dotted histogram (Background IV) is the $J/\psi\rightarrow\gamma\pi^{+}\pi^{-}\pi^{+}\pi^{-}$ MC shape.
V.3 $CP$-violation Asymmetry
The fit function for the angular distribution, based on the squared decay amplitude Petri (2010), is
$F(\varphi)=1+a\cdot\sin^{2}{\varphi}+b\cdot\sin{2\varphi},$ (9)
where $1+a\cdot\sin^{2}{\varphi}$ is a dominant contribution from the magnetic term, and $b\cdot\sin{2\varphi}$ is from the $CP$-violating interference term. Here, $\varphi$ is the angle between the decay planes of the $\pi^{+}\pi^{-}$ and $e^{+}e^{-}$ systems. Then the asymmetry parameter $\mathcal{A}_{CP}$ is defined as
$\mathcal{A}_{CP}$ $=\frac{1}{\Gamma}\int^{2\pi}_{0}\frac{d\Gamma}{d\varphi}\,\rm{sign}(\sin{2\varphi})\,d\varphi$ (10)
[…]
$=\frac{4b}{(2+a)\pi},$
where $\mathrm{sign}(x)$ is the sign function, $\mathrm{sign}(x)=x/|x|$.
The background treatment follows the same method as the TFF measurement (Section V.2), and the background yields are listed in Table 2. In the fit, the efficiency-corrected $F(\varphi)$ function is convolved with a Gaussian function to account for the $\varphi$ resolution. The fit result of $\varphi$ is shown in Fig. 6; the $\chi^{2}/ndf$ is $79.8/97$. With the fitted parameters $a=-0.696\pm 0.061$ and $b=-0.041\pm 0.048$, $\mathcal{A}_{CP}$ is calculated to be
$\mathcal{A}_{CP}(\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-})=(-4.04\pm 4.69)\%,$ (11)
where the uncertainty is statistical only.
Figure 6: Fit to the distribution of plane angle $\varphi$ for $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ decay mode. The points with error bars represent the data, the blue solid histogram is the total fit result, and the red dashed line represents the function shape of $F(\varphi)$ convolved with a Gaussian function. The light blue dashed histogram (Background I) is the $J/\psi\rightarrow\gamma\eta^{\prime},\eta^{\prime}\rightarrow\pi^{+}\pi^{-}\eta,\eta\rightarrow\gamma e^{+}e^{-}$ MC shape. The green dashed histogram (Background II) is the $J/\psi\rightarrow\gamma\eta,\eta\rightarrow\pi^{+}\pi^{-}\pi^{0}$ MC shape. The purple dashed histogram (Background III) is the $J/\psi\rightarrow\gamma\eta,\eta\rightarrow\gamma\pi^{+}\pi^{-}$ MC shape. The pink dotted histogram (Background IV) is the $J/\psi\rightarrow\gamma\pi^{+}\pi^{-}\pi^{+}\pi^{-}$ MC shape.
VI Systematic uncertainties
Sources of systematic uncertainty are summarized in Table 3, and their corresponding contributions are discussed in detail below.
  • Number of $J/\psi$ events: The number of $J/\psi$ events is determined to be $(10087\pm 44)\times 10^{6}$ from counting the inclusive hadronic events; the uncertainty is 0.4% Ablikim et al. (2022).
  • Branching fraction of $J/\psi\rightarrow\gamma\eta$: The uncertainty of 1.2% on the branching fraction of $J/\psi\rightarrow\gamma\eta$ is taken from the PDG Navas et al. (2024).
  • MC statistics: The systematic uncertainty related to the finite statistics of the MC simulation used to obtain the overall reconstruction efficiency is calculated as $\sqrt{\frac{\varepsilon(1-\varepsilon)}{N}}$, where $\varepsilon$ is the detection efficiency and $N$ is the number of generated MC events of the signal process. The corresponding systematic uncertainty is 0.2%.
  • MDC tracking: The data-MC efficiency difference for pion track-finding is studied using a control sample of $J/\psi\rightarrow\pi^{+}\pi^{-}\pi^{0}$. Since there is no specific decay process available for us to study the tracking of muons in the low momentum region and the muon and pion masses are similar, we use the pion study results for muons as well. For the MDC tracking efficiency of electrons, a mixed sample of $e^{+}e^{-}\rightarrow e^{+}e^{-}\gamma$ at the $J/\psi$ meson mass and $J/\psi\rightarrow e^{+}e^{-}\gamma_{FSR}$ is used, where FSR stands for final-state radiation. The data-MC difference, $\Delta_{\rm{syst.}}$, is extracted as a function of the particle momentum and the cosine of the polar angle. Subsequently, each event in the MC samples is re-weighted by a factor $(1+\Delta_{\rm{syst.}})$. The branching fraction and asymmetry parameter are recalculated with efficiencies determined from the re-weighted MC sample. For the TFF measurement, a re-weighted MC sample is used to calculate the MC integral, and a new set of fit results is obtained by using the fit method as outlined in Section V.2. The differences from the original results are taken as the systematic uncertainties.
  • Photon detection efficiency: The photon detection efficiency in the EMC is studied using a control sample of $e^{+}e^{-}\rightarrow\gamma_{\mathrm{ISR}}\mu^{+}\mu^{-}$, where ISR stands for initial-state radiation. The systematic uncertainty is taken as 0.5%, which is the maximum difference in efficiency between data and MC simulation in both the barrel and end-cap regions.
  • PID: The pion PID efficiency for data agrees within 1.0% with that of the MC simulation in the pion momentum region, as reported in Ablikim et al. (2021). There is no specific decay process available for us to study the PID of muons in the low momentum region. Because the muon mass is similar to the pion mass, 1.0% is taken as the systematic uncertainty for the muon Ablikim et al. (2021). Thus, 4.0% is taken as the systematic uncertainty from PID for the $\eta\rightarrow\pi^{+}\pi^{-}\mu^{+}\mu^{-}$ decay.
  • Kinematic fit: To investigate the systematic uncertainty associated with the kinematic fit, the track helix parameter correction method Ablikim et al. (2013) is used. Half of the difference in the detection efficiencies with and without the helix corrections is taken as the systematic uncertainty.
  • Combined PID and kinematic fit: A clean control sample of $J/\psi\rightarrow\pi^{+}\pi^{-}\pi^{0},\pi^{0}\rightarrow\gamma e^{+}e^{-}$ is used to study the systematic uncertainty due to the requirement of $\chi^{2}_{\rm{sum}}(\pi^{+}\pi^{-}e^{+}e^{-})<50$. This sample included both $\pi^{0}$ Dalitz decay and $\pi^{0}\rightarrow\gamma\gamma$ decay with one photon externally converting to an electron-positron pair. Using the same approach as that used for the tracking efficiency, we perform a two-dimensional correction to the selection efficiency of $\chi^{2}_{\rm{sum}}(\pi^{+}\pi^{-}e^{+}e^{-})<50$ as a function of the momentum of the electron and positron. The differences from the original results are taken as the systematic uncertainties.
  • Photon conversion rejection: The control sample of $J/\psi\rightarrow\pi^{+}\pi^{-}\pi^{0},\pi^{0}\rightarrow\gamma e^{+}e^{-}$ is used to evaluate the systematic uncertainty from the rejection of photon conversions. Using the same approach as that used for the tracking efficiency, we perform a two-dimensional correction to the photon conversion rejection efficiency as a function of the momenta of the electrons and positrons. The differences from the original results are taken as the systematic uncertainties.
  • Signal shape: We use a double Gaussian function instead of the signal MC shape to fit $\pi^{+}\pi^{-}e^{+}e^{-}$ mass spectrum. The fit yields $658\pm 27$ signal events and the systematic uncertainty is determined as the difference from the nominal fit to be $3.2\%$.
  • Generator model: A MC generator based on a theoretical calculation is used to determine the detection efficiency for $\pi^{+}\pi^{-}e^{+}e^{-}$ decays. The dependence of the detection efficiency on the form factor is evaluated by replacing the nominal $m_{V}=775\ \mathrm{MeV}/c^{2}$ with the $m_{V}=749\ \mathrm{MeV}/c^{2}$ measurement as described in Section V.2. The difference of the detection efficiency between original model and model (I) is taken as the uncertainty.
  • Background estimation: The background simulations are based on the amplitude analysis and are relatively precise. Therefore, the systematic uncertainty from background models is neglected. In the fit, the numbers of events for specific backgrounds are fixed according to the branching fractions from the PDG. To estimate the effect of the uncertainties of the used branching fractions, a set of random numbers is generated within the uncertainty range of each branching fraction. Using these random scaling parameters, a series of fits to the invariant mass distributions of $\pi^{+}\pi^{-}e^{+}e^{-}$ are performed. The variance of the determined number of signal events and asymmetry parameters is taken as the systematic uncertainties. In the TFF measurement, we perform alternative fits by changing the branching fraction by $-1.0\sigma$ and $+1.0\sigma$, and the largest difference from the nominal results is taken as the systematic uncertainty.
  • Resolution: To estimate the uncertainty from the resolution, we perform alternative fits by changing the resolution from $-1.0\sigma$ to $+1.0\sigma$ for the $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ decay. The largest difference from the nominal result is taken as the systematic uncertainty on the asymmetry parameter.
  • Width: In the TFF measurement, the width of $\rho(770)$ is taken as a constant. The difference of the fit results between the cases of fixing and allowing the width to float is taken as the systematic uncertainty.
Table 3: The systematic uncertainties for the $\eta\rightarrow\pi^{+}\pi^{-}l^{+}l^{-}$ (with $l=e$ or $\mu$) decays.
Source $l=e$ (%) $l=\mu$ (%)
[…]
Number of $J/\psi$ events $0.4$.. ... $0.4$
$\mathcal{B}(J/\psi\rightarrow\gamma\eta)$ $1.2$.. ... $1.2$
MC statistics $0.2$.. ... $0.2$
MDC tracking $4.1$.. $1.0$ $0.7$
Photon detection $0.5$.. ... $0.5$
PID ... ... ... $4.0$
Kinematic fit ... ... ... $0.5$
Combine PID and kinematic fit $3.1$ $0.9$ $0.2$..
Photon conversion rejection $1.1$ $1.6$ $1.6$..
Signal shape $3.2$.. ... ...
Generator model $0.1$.. ... ...
Background estimation $0.5$ $2.2$ $0.0$..
Resolution ... $1.8$.. ...
Width ... ... $0.3$..
Total $6.3$ $3.4$ $1.9$ $4.3$
VII Upper limit of $\eta\rightarrow\pi^{+}\pi^{-}\mu^{+}\mu^{-}$
No events are found in the $\eta$ mass region ($N_{\rm{sig}}=0$). The number of remaining background events from all $J/\psi$ events is normalized to be $N_{\rm{bkg}}=0$. The detection efficiency is $\varepsilon_{\rm{sig}}=(45.65\pm 0.08)\%$ and the total systematic uncertainty is $\Delta_{\rm{sys.}}=4.3\%$. The TROLKE method is useful in situations with low statistical significance, and the limit calculations make use of the profile likelihood method Lundberg et al. (2010). After inserting all the above numbers in the TROLKE program, the upper limit of signal events is obtained to be $N^{\mathrm{UP}}=4.4$ at the 90% confidence level (C.L.). Thus, the upper limit on the branching fraction at the 90% C.L. is determined as
$\mathcal{B}<\frac{N^{\mathrm{UP}}}{N_{J/\psi}\cdot\mathcal{B}(J/\psi\rightarrow\gamma\eta)}=4.0\times 10^{-7}.$ (12)
VIII Upper limits for ALPs in $\eta\rightarrow\pi^{+}\pi^{-}a,a\rightarrow e^{+}e^{-}$
To study the unknown ALPs decaying into $e^{+}e^{-}$, we perform 40 simulations under the hypothesis $a\rightarrow e^{+}e^{-}$ by varying masses of $a$ in steps of $5$ MeV/$c^{2}$ from 0 to 200 MeV/$c^{2}$.
The $a$ signal shape is obtained through MC simulation with an assumption of a negligible width. The treatment of the background is the same as that of the TFF measurement in Section V.2, and the normalized background yields are listed in Table 2. Then 40 unbinned maximum likelihood fits to the $e^{+}e^{-}$ invariant mass spectrum are performed.
We consider possible sources for multiplicative systematic uncertainties of the upper limits for ALPs, such as number of $J/\psi$ events (0.4%), branching fraction of $J/\psi\rightarrow\gamma\eta$ (1.2%), MC statistics (0.1%), MDC tracking ($2.6\%-5.0\%$), photon detection efficiency (0.5%), combined PID and kinematic fit (3.1%) and photon conversion veto (1.1%). The total multiplicative systematic uncertainties for each hypothesis of $a$ mass are $(4.4-6.1)\%$. The additive systematic uncertainties are considered by alternative fit ranges and Background estimation. The maximum signal yield among the different fit scenarios is adopted as its upper limit.
Since no evident $a$ signals are seen in the $M(e^{+}e^{-})$ distribution, we compute the upper limits on the relative branching fraction, $R^{\mathrm{UP}}=\frac{\mathcal{B}(\eta\rightarrow\pi^{+}\pi^{-}a)\cdot\mathcal{B}(a\rightarrow e^{+}e^{-})}{\mathcal{B}(\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-})}$, at the 90% C.L. as a function of $M(a)$. The upper limits on the number of $a$ signal events at the 90% C.L. are obtained according to the Bayesian method Zhu (2008) by smearing the likelihood curve using a Gaussian function with a width of the systematic uncertainty as
$L^{\prime}(N)=\int^{1}_{0}L\left(\frac{S}{\hat{S}}N\right)\mathrm{exp}\left[-\frac{(S-\hat{S})^{2}}{2\sigma_{S}^{2}}\right]\mathrm{d}S,$ (13)
where $L$ and $L^{\prime}(N)$ are the likelihood curves before and after taking into account the systematic uncertainties; $\hat{S}$ is the nominal efficiency and $\sigma_{S}$ is its systematic uncertainty. As shown in Fig. 7, the combined limits on the relative branching fractions are established at the level of $(3.0-53.0)\times 10^{-3}$, and the significance for each case less than $0.5\sigma$.
Figure 7: Upper limits on the relative width at the 90% C.L. for different $a$ masses.
IX SUMMARY
With a sample of $(10087\pm 44)\times 10^{6}$ $J/\psi$ events, the decays of $\eta\rightarrow\pi^{+}\pi^{-}l^{+}l^{-}$ ($l=e$ or $\mu$) are studied. For the decay $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$, the branching fraction is determined to be $\mathcal{B}(\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-})=(3.07\pm 0.12_{\rm{stat.}}\pm 0.19_{\rm{syst.}})\times 10^{-4}$, which is in good agreement with theoretical predictions Petri (2010); Borasoy and Nissler (2007) and previous measurements Berlowski et al. (2008); Akhmetshin et al. (2001); Ambrosino et al. (2009), as shown in Fig. 8.
Figure 8: The branching fraction of $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ from different theoretical predictions Petri (2010); Borasoy and Nissler (2007) (blue squares), other experiments Berlowski et al. (2008); Akhmetshin et al. (2001); Ambrosino et al. (2009) (red triangles) and this measurement (green dot).
For the decay $\eta\rightarrow\pi^{+}\pi^{-}\mu^{+}\mu^{-}$, no signal events are observed. The upper limits on the branching fraction are determined to be $\mathcal{B}(\eta\rightarrow\pi^{+}\pi^{-}\mu^{+}\mu^{-})<4.0\times 10^{-7}$ at the 90% C.L., which is improved by 3 orders of magnitude compared to the PDG value $3.6\times 10^{-4}$.
Furthermore, the TFF is extracted from the invariant decay amplitude of the reaction $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$. An analysis on previous experimental data obtained the values of parameters $c_{1}-c_{2}$ and $c_{3}$ that are approximately equal to $1$ Benayoun et al. (2010) and consistent with Model I (hidden gauge), we thus report our results based on this model. We obtain $m_{V}=(749\pm 54_{\rm{stat.}}\pm 14_{\rm{syst.}})$ $\mathrm{MeV}/c^{2}$ for the hidden gauge model; a larger $\eta$ data sample is needed to improve precision.
Additionally, the $CP$-violation asymmetry is determined to be $\mathcal{A}_{CP}(\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-})=(-4.04\pm 4.69_{\rm{stat.}}\pm 0.14_{\rm{syst.}})\%$, which implies no $CP$-violation under the present statistics.
Finally, ALPs are searched for via the decay $\eta\rightarrow\pi^{+}\pi^{-}a,a\rightarrow e^{+}e^{-}$, and the 90% C.L. upper limits on the branching fraction relative to that of $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ are presented as a function of the ALP mass within $5-200\ \mathrm{MeV}/c^{2}$, as shown in Fig. 7. The significance for each case is less than $0.5\sigma$.
Acknowledgements.
The BESIII Collaboration thanks the staff of BEPCII and the IHEP computing center for their strong support. This work is supported in part by National Key R&D Program of China under Contracts Nos. 2020YFA0406300, 2020YFA0406400, 2023YFA1606000; National Natural Science Foundation of China (NSFC) under Contracts Nos. 11635010, 11735014, 11935015, 11935016, 11935018, 12025502, 12035009, 12035013, 12061131003, 12192260, 12192261, 12192262, 12192263, 12192264, 12192265, 12221005, 12225509, 12235017, 12475089, 12361141819; the Chinese Academy of Sciences (CAS) Large-Scale Scientific Facility Program; the CAS Center for Excellence in Particle Physics (CCEPP); Joint Large-Scale Scientific Facility Funds of the NSFC and CAS under Contract No. U1832207; 100 Talents Program of CAS; The Institute of Nuclear and Particle Physics (INPAC) and Shanghai Key Laboratory for Particle Physics and Cosmology; German Research Foundation DFG under Contracts Nos. 455635585, FOR5327, GRK 2149; Istituto Nazionale di Fisica Nucleare, Italy; Ministry of Development of Turkey under Contract No. DPT2006K-120470; National Research Foundation of Korea under Contract No. NRF-2022R1A2C1092335; National Science and Technology fund of Mongolia; National Science Research and Innovation Fund (NSRF) via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation of Thailand under Contract No. B16F640076; Polish National Science Centre under Contract No. 2019/35/O/ST2/02907; The Swedish Research Council; U. S. Department of Energy under Contract No. DE-FG02-05ER41374.